# Learning Massive-scale Partial Correlation Networks in Clinical Multi-omics Studies with HP-ACCORD

### v2

@@ Abstract
Graphical model estimation from modern multi-omics data requires a balance between statistical estimation performance and computational scalability. We introduce a novel pseudolikelihood-based graphical model framework that reparameterizes the target precision matrix while preserving sparsity pattern and estimates it by minimizing an $\ell_{1}$-penalized empirical risk based on a new loss function. The proposed estimator maintains estimation and selection consistency in various metrics under high-dimensional assumptions. The associated optimization problem allows for a provably fast computation algorithm using a novel operator-splitting approach and communication-avoiding distributed matrix multiplication. A high-performance computing implementation of our framework was tested in simulated data with up to one million variables demonstrating complex dependency structures akin to biological networks. Leveraging this scalability, we estimated partial correlation network from a dual-omic liver cancer data set. The co-expression network estimated from the ultrahigh-dimensional data showed superior specificity in prioritizing key transcription factors and co-activators by excluding the impact of epigenomic regulation, demonstrating the value of computational scalability in multi-omic data analysis.
@@ 1 Introduction
Omics technologies are essential discovery tools for the analysis of complex biological samples such as tissues and biofluids. Individual omics modalities can generate high-dimensional data (large $p$) for characterizing complex biological samples in a high-throughput manner (large $n\ll p$), and recently for deconvoluting molecular profiles of isolated single cells or spatially connected micrometer-sized regions over tissue sections [37]. Given the unprecedented opportunity to assemble multi-omic data as fine-resolution descriptors of gene expression regulation, cellular metabolism and signal transduction in a biological system, clinical and basic research studies routinely employ a combination of omics platforms to characterize joint variation in the genome, epigenome, transcriptome, proteome and metabolome [12, 16].
@@
In biomedical applications, the method of choice for multi-omic data interpretation is to map experimental data over well-characterized biological networks, where the data features are linked by a directed graph representing gene expression regulatory network, metabolic pathways or signaling cascades, or an undirected graph such as protein-6protein interaction network or gene co-expression network. Different networks facilitate human interpretation of molecular interactions in their respective biological context. However, a downside of this approach is that the network information biologically relevant to a given study may be unavailable in the literature, or the network information can be incomplete and thus fail to cover the molecules measured in a study. In other cases, the biological network of interest may be adaptive, and it is therefore susceptible to rewiring in specific conditions, rendering the available static network information inapplicable. For these reasons, it is often necessary to complement existing networks by estimating new ones directly from data so that a realistic snapshot of the relational structure among the variables is reflected in the downstream analysis or interpretation of the results.
@@
In omics-scale analysis, biological network is typically represented as an adjacency matrix of data features based on marginal pairwise correlations. In the popular weighted gene co-expression network analysis [WGCNA, 41], for example, Pearson correlation-based network modules are captured first and the initial network is refined by topological analysis. Despite the success of WGCNA and other extensions in identifying gene communities from omics data, the adjacency of two molecular features as determined by marginal correlations can be a byproduct of shared regulatory factors rather than direct biological interaction. Hence it is natural to estimate the co-expression network underlying a gene expression dataset with a graph representing the conditional dependencies, e.g., partial correlations. The same logic extends to the inference of other types of biological networks from multi-omic data such as metabolic pathways and signaling cascades.
@@
It is well known that the precision matrix, or inverse covariance matrix, of a multivariate distribution encodes partial correlations. The most popular high-dimensional precision matrix estimation approach is the graphical lasso [40, 11], which maximizes a multivariate Gaussian likelihood with an $\ell_{1}$-penalty. However, at the scale of modern multi-omics datasets, it is computationally infeasible to estimate desired networks using this strategy mainly due to the computational bottleneck, which is not easily resolved by using a more powerful computer. As we demonstrate later, our application data features close to 300,000 variables. On two high-performance computing (HPC) environments, the most scalable implementation of graphical lasso, BigQUIC [14], could complete the computation only for the trivial cases in which the computed precision matrix estimates were diagonal. Similarly, a fast implementation [30] of CLIME [4], another popular methodology, was also not applicable on a HPC machine with 192GB of memory when the dataset dimension exceeded 30,000. In the next section, we discuss in detail that the inherent nature of the optimization problems from these popular methods makes it difficult to scale up to modern omics-scale problems.
@@
Feature screening procedures, which reduces the number of variables before the network estimation step, have been proposed as a remedy to address the lack of scalability of methods mentioned above [24, 42]. Although feature screening could be suitable when a moderate dimensionality reduction is necessary, an order of magnitude reduction (300,000 or more to 30,000 or fewer features) precludes functionally important molecules from being considered for network estimation for computational tractability rather than a biologically motivated rationale. It is also difficult to discern essential variables and dispensable ones a priori solely based on numerical criteria chosen by an analyst, hence there is a clear merit in estimating the model with all possible input variables included from the start.
@@
In summary, a statistical framework for partial correlation graph estimation scalable to contemporary multi-omics studies is currently lacking. To address this gap, we propose a new estimation framework called ACCORD that aims to strike a balance between computational scalability and statistical performance. In ACCORD, target precision matrix is reparameterized and estimated by minimizing the $\ell_{1}$-penalized empirical risk based on a new loss function. The associated optimization problem allows for a provably fast computation algorithm that is massively scalable, which is realized by a novel operator splitting algorithm and communication-avoiding distributed matrix multiplication. We then show that the ACCORD estimator possesses estimation consistency in $\ell_{1}$ and $\ell_{2}$ norms under standard high-dimensional assumptions and selection consistency under an irrepresentability condition. In this way, we balance between statistical performance and scalability of computation in massively large-scale settings.
We demonstrate that the HPC implementation of ACCORD, termed HP-ACCORD, scales well to handle dimensions up to one million. Leveraging on the scalability, we estimate a partial correlation network in the multi-omic liver cancer dataset from The Cancer Genome Atlas (TCGA). Using a combination of epigenomic and transcriptomic data sets with 285,358 variables in total, we successfully recapitulated local and global correlation structures of the variables within the same omics platform and identified a network of co-transcribed genes and DNA methylation events in upstream regulatory regions. More importantly, we show that the graph estimated from the ultrahigh-dimensional dual-omic data enabled us to identify bona fide transcription factors driving the co-expression network with greater specificity than and equivalent sensitivity to the alternative graph estimated from the transcriptomic data with 15,598 variables only. The analysis clearly highlights the merit of performing graph estimation in the whole feature space.
The organization of this paper is as follows. Section 2 explains the challenges that existing methods face when the scale of the problem becomes massively large. Section 3 introduces the ACCORD framework. In Section 4 we study the statistical properties of the ACCORD estimator. Section 5 is devoted to numerical experiments demonstrating the performance of the estimator and scalability of the algorithm. In Section 6, we showcase HP-ACCORD through graph estimation in a liver cancer data set with 285,358 multi-omic variables. The paper is concluded in Section 7.
@@ 2 Challenges
In this section we first describe the computational hurdles that current techniques for high-dimensional precision matrix estimation grapple with when dealing with extremely large scales.
Graphical lasso solves the convex optimization problem $\min_{\boldsymbol{\Theta}\in\mathbb{S}^{p}}\{-\log\det\boldsymbol{\Theta}+\operatorname{\bf tr}(\boldsymbol{S}\boldsymbol{\Theta})+\lambda\|\boldsymbol{\Theta}\|_{1}\}$, $\lambda>0$, in order to estimate the precision matrix $\boldsymbol{\Theta}^{*}$ of a $p$-variate zero-mean distribution. Here $\mathbb{S}^{p}$ is the space of $p\times p$ symmetric matrices, $\boldsymbol{S}=(1/n)\boldsymbol{X}^{T}\boldsymbol{X}$ is the sample covariance matrix, where $\boldsymbol{X}=[X_{1},X_{2},\cdots,X_{n}]^{T}$ with $X_{i}\in\mathbb{R}^{p}$ is the centered data matrix, and $\|\cdot\|_{1}$ is the vector $\ell_{1}$ norm. The estimate is fully characterized by the Karush-Kuhn-Tucker (KKT) optimality condition
@@
where $\partial\|\boldsymbol{\Theta}\|_{1}$ denotes the subdifferential of the convex function $\boldsymbol{X}\mapsto\|\boldsymbol{X}\|_{1}$ at $\boldsymbol{\Theta}$. The vast literature on the algorithms for graphical lasso, e.g., [8, 11, 23, 13, 15], essentially reduces to how to solve the KKT equation (2.1) iteratively.
@@
The computational culprit in (2.1) is the inverse of the $p\times p$ matrix variable $\boldsymbol{\Theta}$, which has to be computed every iteration. Inverting a $p\times p$ matrix using direct methods (e.g., Cholesky decomposition) costs $O(p^{3})$ arithmetic operations in general, which becomes prohibitive if $p$ is at the omics scale of a few hundreds of thousands. Storage requirements worsen the situation. The inverse $\boldsymbol{\Theta}^{-1}$ is in general not sparse, requiring $O(p^{2})$ memory space. For example if $p=300,000$, then roughly 700GB of memory is needed, which calls for distributed computation. Being inherently sequential, however, matrix inversion is difficult to parallelize or distribute, meaning that direct methods are not so much scalable even with high-performance computing (HPC) systems with distributed memory.
@@
Consequently, most existing algorithms resort to employing an (inner) iterative method to compute $\boldsymbol{\Theta}^{-1}$. For instance, Friedman et al. [11] invoke a $p$-dimensional lasso regression solver $p$ times every (outer) iteration to compute one column of $\boldsymbol{\Theta}^{-1}$ at a time, which becomes prohibitively expensive in omics-scale problems. BigQUIC [14] updates $\boldsymbol{\Theta}$ by block coordinate descent and then computes $\boldsymbol{\Theta}^{-1}$ column by column, by storing only a small fraction of the $p$ columns of $\boldsymbol{\Theta}^{-1}$ in a cache and recomputing the missing columns on demand; the recomputation solves a fraction of the $p$ linear equations defining the inverse ($\boldsymbol{\Theta}\boldsymbol{X}=\boldsymbol{I}_{p}$) by using conjugate gradient. Since this strategy is only successful if the cache miss rate is low, BigQUIC partitions $\boldsymbol{\Theta}$ into a block matrix and permutes blocks to minimize the number of the off-block-diagonal elements. Therefore, the scalability of BigQUIC is limited by the maximum degree of the underlying graph. For more complex graph structures of interest, e.g., those that arise in multi-omics studies, computable dimensions are practically less than $100,000$; see §5.
@@
CLIME minimizes $\|\boldsymbol{\Theta}\|_{1}$ subject to the constraint $\|\boldsymbol{S}\boldsymbol{\Theta}-\boldsymbol{I}_{p}\|_{\infty}\leq\lambda$ over $\mathbb{R}^{p\times p}$, where $\|\cdot\|_{\infty}$ is the vector $\ell_{\infty}$ norm. The optimization problem of CLIME can be decomposed into $p$ independent, $p$-dimensional linear programming (LP) problems, each solving a column of $\boldsymbol{\Theta}$. While LP is the most extensively studied convex optimization problem, commercial LP solvers like Gurobi or Cplex, which utilize interior-point or simplex methods, struggle with scalability when handling hundreds of thousands of variables [25]. Exacerbating the situation, for omics-scale CLIME we are required to solve hundreds of thousands of large-scale LP problems. This remains an almost unattainable task despite the independence. As a result, fastclime [30, 31], which could not handle more than 30,000 variables in our problem instance, remains one of the most scalable implementations publicly available.
@@
As the most scalable pseudolikelihood-based method up to date, CONCORD [18] solves the $\ell_{1}$-penalized minimization problem
@@
Compared with that for graphical lasso (2.1), KKT condition (2.3) only involves the inverse of the diagonal matrix $\boldsymbol{\Theta}_{D}$, which is trivial to compute. While the second term in (2.3) costs two $p\times p$ matrix-matrix multiplications with $O(p^{3})$ operations, they are much easier to distribute and parallelize than matrix inversion. The sparsity of the optimization variable $\boldsymbol{\Theta}$ imposed by the $\ell_{1}$ penalty further reduces the complexity of the multiplication. Koanantakool et al. [20] leverage these features to achieve massive scalability in distributed memory HPC systems. This HP-CONCORD implementation can handle 320,000 dimensional data drawn from a random graph structure resembling multi-omic networks.
Despite its computational success, a major drawback of CONCORD is its lack of statistical consistency in estimating $\boldsymbol{\Theta}^{*}$, which is important in downstream tasks in multi-omic analyses as it carries information on the strength of network edges. The objective function in (2.2) defines a loss function $L(\boldsymbol{\Theta},x)=-\log\det\boldsymbol{\Theta}_{D}+\frac{1}{2}\operatorname{\bf tr}(x^{T}\boldsymbol{\Theta}^{2}x)$ for a sample $x\in\mathbb{R}^{p}$. The population risk $R(\boldsymbol{\Theta})=\operatorname{\bf E}L(\boldsymbol{\Theta},X)$ is minimized by a $\boldsymbol{\Theta}$ satisfying the equation $\boldsymbol{\Theta}_{D}^{-1}-\frac{1}{2}(\boldsymbol{\Theta}\boldsymbol{\Sigma}^{*}+\boldsymbol{\Sigma}^{*}\boldsymbol{\Theta})=\mathbf{0}$, where $\boldsymbol{\Sigma}^{*}={\boldsymbol{\Theta}^{*}}^{-1}$ is the covariance matrix of random vector $X$. Unfortunately, the true precision matrix $\boldsymbol{\Theta}^{*}$ does not minimize the risk unless $\boldsymbol{\Theta}_{D}$ equals the identity. This is a very restrictive assumption indeed, and likely does not hold in many applications. Consequently, the currently available result on the consistency of the CONCORD estimator [Theorem 2, 18] requires an accurate estimator $\widehat{\boldsymbol{\Theta}_{D}}$ of $\boldsymbol{\Theta}^{*}_{D}$ with a rate $\max_{i=1,\dotsc,p}|\widehat{\boldsymbol{\Theta}_{D}}_{,ii}-\boldsymbol{\Theta}^{*}_{D,ii}|=O_{P}(\sqrt{\log n/n})$. Similarly, SPACE also requires a separate estimator of $\boldsymbol{\Theta}_{D}^{*}$ with the same rate. Such a separate estimator is difficult to find in practice.
@@
There is also an issue with the optimization method employed by HP-CONCORD. This CONCORD-ISTA algorithm [29] possesses an $O(1/t)$ rate of convergence, where $t$ is the number of iterations, and the convergence is measured in terms of the objective function value. In addition to the slow, sublinear convergence of the objective function, the variable iterate $\{\boldsymbol{\Theta}^{(t)}\}$ may converge to the minimizer $\hat{\boldsymbol{\Theta}}$ of (2.2) in an arbitrarily slow rate [2]. Since the algorithm has to be terminated within a finite number of iterations, statistical error of $\boldsymbol{\Theta}^{(t)}$ (say $\|\boldsymbol{\Theta}^{(t)}-\boldsymbol{\Theta}^{*}\|$) may remain quite large compared to that of $\hat{\boldsymbol{\Theta}}$ (say $\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{*}\|)$ even after a large number of iterations.

### v3

@@ Abstract
Graphical model estimation from multi-omics data requires a balance between statistical estimation performance and computational scalability. We introduce a novel pseudolikelihood-based graphical model framework that reparameterizes the target precision matrix while preserving the sparsity pattern and estimates it by minimizing an $\ell_{1}$-penalized empirical risk based on a new loss function. The proposed estimator maintains estimation and selection consistency in various metrics under high-dimensional assumptions. The associated optimization problem allows for a provably fast computation algorithm using a novel operator-splitting approach and communication-avoiding distributed matrix multiplication. A high-performance computing implementation of our framework was tested using simulated data with up to one million variables, demonstrating complex dependency structures similar to those found in biological networks. Leveraging this scalability, we estimated a partial correlation network from a dual-omic liver cancer data set. The co-expression network estimated from the ultrahigh-dimensional data demonstrated superior specificity in prioritizing key transcription factors and co-activators by excluding the impact of epigenetic regulation, thereby highlighting the value of computational scalability in multi-omic data analysis.
@@ 1 Introduction
Omics-scale technologies are valuable tools for surveying molecular features in biological samples such as tissues and biofluids in an unbiased manner. Individual omics modalities can generate high-dimensional data (large $p$) to characterize complex biological samples ($n\ll p$) and, more recently, isolated single cells or spatially connected micrometer-sized regions of tissue sections ($n\approx p$) [40]. Given the unprecedented opportunity to assemble multi-omic data as fine-resolution descriptors of gene expression regulation, cellular metabolism and signal transduction in a biological system, research studies in molecular medicine routinely employ a combination of omics platforms to study joint variation in the genome, epigenome, transcriptome, proteome and metabolome [13, 17].
@@
In biomedical applications, it is customary to map multi-omic data to well-characterized biological pathways or networks for the integrative interpretation of results. In network-based analysis, data features are linked through a directed graph representing gene expression regulatory network, metabolic reaction pathways or signaling cascades, or an undirected graph such as protein-protein interaction network or gene co-expression network. Different networks facilitate the human interpretation of molecular interactions within their respective biological contexts. However, a downside of this approach is the requirement that network information relevant to a given study be available in the literature, covering all molecules measured in a study. This assumption is rarely satisfied in practice. In other cases, the biological network of interest may be adaptive, and it is therefore susceptible to rewiring under a specific condition, rendering the available static network information inapplicable. For these reasons, it is often necessary to complement existing networks by directly estimating new ones from data, ensuring that a realistic snapshot of the relational structure among variables is reflected in the downstream analysis or interpretation of the results.
@@
In omics-scale analysis, the biological network is typically represented as an adjacency matrix of data features based on marginal pairwise correlations. In weighted gene co-expression network analysis [WGCNA, 44], for example, Pearson correlation-based network modules are captured first, and the initial network is refined by topological analysis. Despite the success of WGCNA and other extensions in identifying gene communities from omics data, the adjacency of two molecular features as determined by marginal correlations may be a by-product of shared regulatory factors rather than a direct biological interaction. Hence, it is natural to estimate the co-expression network underlying a gene expression dataset by using a graph that represents the conditional dependencies, e.g., partial correlations. The same logic extends to the inference of other types of biological networks from multi-omic data, such as metabolic pathways and signaling cascades.
@@
It is well known that the precision matrix, or inverse covariance matrix, of a multivariate distribution encodes partial correlations. The most popular high-dimensional precision matrix estimation approach is the graphical lasso [43, 11], which maximizes a multivariate Gaussian likelihood with an $\ell_{1}$-penalty. However, at the scale of modern multi-omics datasets, it is computationally infeasible to estimate desired networks using this strategy, mainly due to the computational bottleneck that is not easily resolved by using a more powerful computer. As we demonstrate later, our application data features close to 300,000 variables. In two high-performance computing (HPC) environments, the most scalable implementation of graphical lasso, BigQUIC [15], a scaled version of the quadratic approximation for sparse inverse covariance learning (QUIC) algorithm [14], could complete the computation only for the trivial cases in which the computed precision matrix estimates were diagonal. Similarly, a fast implementation [30] of the constrained $\ell_{1}$-minimization for inverse matrix estimation [CLIME, 4], another popular methodology, was also not applicable on a HPC machine with 192GB of memory when the data set dimension exceeded 30,000. In the next section, we discuss in detail the inherent nature of the optimization problems from these popular methods that makes them difficult to scale up to modern omics-scale problems.
@@
Feature screening procedures, which reduce the number of variables before the network estimation step, have been proposed as a remedy to address the lack of scalability of methods mentioned above [24, 47]. Although feature screening could be suitable when a moderate dimensionality reduction is necessary, an order of magnitude reduction (300,000 or more to 30,000 or fewer features) precludes functionally important molecules from being considered for network estimation for computational tractability rather than a biologically motivated rationale. It is also difficult to discern essential variables and dispensable ones a priori solely based on numerical criteria chosen by an analyst; therefore, there is a clear merit in estimating the model with all possible input variables included from the start.
@@
In summary, a statistical framework for partial correlation graph estimation, scalable to contemporary multi-omics studies, is currently lacking. To address this gap, we propose a new estimation framework called ACCORD that aims to strike a balance between computational scalability and statistical performance. In ACCORD, target precision matrix is reparameterized and estimated by minimizing the $\ell_{1}$-penalized empirical risk based on a new loss function. The associated optimization problem enables a massively scalable and provably fast computation algorithm, achieved through a novel operator splitting algorithm and communication-avoiding distributed matrix multiplication. We then show that the ACCORD estimator has estimation consistency in $\ell_{1}$ and $\ell_{2}$ norms under standard high-dimensional assumptions and selection consistency under an irrepresentability condition. In this way, we strike a balance between statistical performance and computational scalability in massively large-scale settings.
We demonstrate that the HPC implementation of ACCORD, termed HP-ACCORD, scales well to handle dimensions up to one million. Leveraging on the scalability, we estimate a partial correlation network in the multi-omic liver cancer dataset from The Cancer Genome Atlas (TCGA). Using a combination of epigenomic and transcriptomic data sets with a total of 285,358 variables, we successfully recapitulated the local and global correlation structures of the variables within the same omics platform and identified a network of co-transcribed genes and DNA methylation events in upstream regulatory regions. More importantly, we show that the graph estimated from the ultrahigh-dimensional dual-omic data enabled us to identify bona fide transcription factors driving the co-expression network with greater specificity than and equivalent sensitivity to the alternative graph estimated from the transcriptomic data with 15,598 variables only. The analysis clearly highlights the merit of performing graph estimation in the whole feature space.
The organization of this paper is as follows. Section 2 explains the challenges that existing methods face when the scale of the problem becomes massively large. Section 3 introduces the ACCORD framework. In Section 4 we study the statistical properties of the ACCORD estimator. Section 5 is devoted to numerical experiments demonstrating the performance of the estimator and the scalability of the algorithm. In Section 6, we showcase HP-ACCORD through graph estimation in a liver cancer data set with 285,358 multi-omic variables. The paper is concluded in Section 7.
@@ 2 Challenges
In this section, we first describe the computational hurdles that current techniques for high-dimensional precision matrix estimation grapple with when dealing with extremely large scales.
Graphical lasso solves the convex optimization problem $\min_{\boldsymbol{\Theta}\in\mathbb{S}^{p}}\{-\log\det\boldsymbol{\Theta}+\operatorname{\bf tr}(\boldsymbol{S}\boldsymbol{\Theta})+\lambda\|\boldsymbol{\Theta}\|_{1}\}$, $\lambda>0$, in order to estimate the precision matrix $\boldsymbol{\Theta}^{*}$ of a $p$-variate zero-mean distribution. Here $\mathbb{S}^{p}$ is the space of $p\times p$ symmetric matrices, $\boldsymbol{S}=(1/n)\boldsymbol{X}^{T}\boldsymbol{X}$ is the sample covariance matrix, where $\boldsymbol{X}=[X_{1},X_{2},\cdots,X_{n}]^{T}$ with $X_{i}\in\mathbb{R}^{p}$ is the centered data matrix, and $\|\cdot\|_{1}$ is the vector $\ell_{1}$ norm. The smooth part of the objective is a simplified form of the negative log-likelihood of a zero-mean normal distribution; the $\ell_{1}$ penalty promotes the sparsity of the estimate.The estimate is fully characterized by the Karush-Kuhn-Tucker (KKT) optimality condition
@@
where $\partial\|\boldsymbol{\Theta}\|_{1}$ denotes the subdifferential of the convex function $\boldsymbol{X}\mapsto\|\boldsymbol{X}\|_{1}$ at $\boldsymbol{\Theta}$. The vast literature on the algorithms for graphical lasso, e.g., d’Aspremont et al. [8], Friedman et al. [11], Li et al. [23], Hsieh et al. [14, 16], essentially reduces to how to solve the KKT equation (2.1) iteratively.
@@
The computational culprit in (2.1) is the inverse of the $p\times p$ matrix variable $\boldsymbol{\Theta}$, which has to be computed every iteration. Inverting a $p\times p$ matrix using direct methods (e.g., Cholesky decomposition) costs $O(p^{3})$ arithmetic operations in general, which becomes prohibitive if $p$ is at the omics scale of a few hundred thousand. Storage requirements worsen the situation. The inverse $\boldsymbol{\Theta}^{-1}$ is in general not sparse, requiring $\Omega(p^{2})$ memory space. For example, if $p=300,000$, then roughly 700GB of memory is needed, which calls for distributed computation. Being inherently sequential, however, matrix inversion is difficult to parallelize or distribute, meaning that direct methods are not scalable even with high-performance computing (HPC) systems that have distributed memory.
@@
Consequently, most existing algorithms resort to employing an (inner) iterative method to compute $\boldsymbol{\Theta}^{-1}$. For instance, Friedman et al. [11] invokes a $p$-dimensional lasso regression solver $p$ times every (outer) iteration to compute one column of $\boldsymbol{\Theta}^{-1}$ at a time, which becomes prohibitively expensive in omics-scale problems. BigQUIC [15] updates $\boldsymbol{\Theta}$ by block coordinate descent and then computes $\boldsymbol{\Theta}^{-1}$ column by column, by storing only a small fraction of the $p$ columns of $\boldsymbol{\Theta}^{-1}$ in a cache and recomputing the missing columns on demand; the recomputation solves a fraction of the $p$ linear equations defining the inverse ($\boldsymbol{\Theta}\boldsymbol{X}=\boldsymbol{I}_{p}$) by using conjugate gradient. Since this strategy is only successful if the cache miss rate is low, BigQUIC partitions $\boldsymbol{\Theta}$ into a block matrix and permutes blocks to minimize the number of off-block-diagonal elements. Therefore, the scalability of BigQUIC is limited by the maximum degree of the underlying graph. For more complex graph structures of interest, e.g., those that arise in multi-omics studies, computable dimensions are practically less than $100,000$; see §5.
@@
CLIME minimizes $\|\boldsymbol{\Theta}\|_{1}$ subject to the constraint $\|\boldsymbol{S}\boldsymbol{\Theta}-\boldsymbol{I}_{p}\|_{\infty}\leq\lambda$ over $\mathbb{R}^{p\times p}$, where $\|\cdot\|_{\infty}$ is the vector $\ell_{\infty}$ norm. This constraint is a margin-allowed version of the estimating equation $\boldsymbol{S}\boldsymbol{\Theta}=\boldsymbol{I}_{p}$ in case $\boldsymbol{S}$ is singular.The optimization problem of CLIME can be decomposed into $p$ independent, $p$-dimensional linear programming (LP) problems, each solving a column of $\boldsymbol{\Theta}$. While LP is the most extensively studied convex optimization problem, commercial LP solvers like Gurobi or Cplex, which utilize interior-point or simplex methods, struggle with scalability when handling hundreds of thousands of variables [25]. Exacerbating the situation, for omics-scale CLIME we are required to solve hundreds of thousands of large-scale LP problems. This remains an almost unattainable task despite the independence. As a result, fastclime [30, 31], which could not handle more than 30,000 variables in our problem instance, remains one of the most scalable implementations publicly available.
@@
As the most scalable pseudolikelihood-based method to date, the convex correlation selection method [CONCORD, 19] solves the $\ell_{1}$-penalized minimization problem
@@
Compared with that for graphical lasso (2.1), KKT condition (2.3) only involves the inverse of the diagonal matrix $\boldsymbol{\Theta}_{D}$, which is trivial to compute. While the second term in (2.3) costs two $p\times p$ matrix-matrix multiplications with $O(p^{3})$ operations, they are much easier to distribute and parallelize than matrix inversion. The sparsity of the optimization variable $\boldsymbol{\Theta}$ imposed by the $\ell_{1}$ penalty further reduces the complexity of the multiplication. Koanantakool et al. [21] leverage these features to achieve massive scalability in distributed memory HPC systems. This HP-CONCORD implementation can handle 320,000-dimensional data drawn from a random graph structure resembling multi-omic networks.
The CONCORD method, while innovative, has significant room for improvement on both computational and statistical fronts. Statistically, the CONCORD approach is not consistent in estimating $\boldsymbol{\Theta}^{*}$, which provides key information on the strength of the edges in the underlying graph. The objective function in (2.2) defines a loss function $L(\boldsymbol{\Theta},x)=-\log\det\boldsymbol{\Theta}_{D}+\frac{1}{2}\operatorname{\bf tr}(x^{T}\boldsymbol{\Theta}^{2}x)$ for a sample $x\in\mathbb{R}^{p}$. The population risk $R(\boldsymbol{\Theta})=\operatorname{\bf E}L(\boldsymbol{\Theta},X)$ is minimized by a $\boldsymbol{\Theta}$ satisfying equation
$\boldsymbol{\Theta}_{D}^{-1}-(1/2)(\boldsymbol{\Theta}\boldsymbol{\Sigma}^{*}+\boldsymbol{\Sigma}^{*}\boldsymbol{\Theta})=\mathbf{0},$ (2.4)
@@
where $\boldsymbol{\Sigma}^{*}={\boldsymbol{\Theta}^{*}}^{-1}$ is the covariance matrix of the random vector $X$. However, the true precision matrix $\boldsymbol{\Theta}^{*}$ does not minimize the risk unless $\boldsymbol{\Theta}_{D}^{*}$ is equal to the identity—a restrictive assumption that is unlikely to hold in practice. The currently available result on the consistency of the CONCORD estimator [Theorem 2, 19] requires an accurate estimator $\widehat{\boldsymbol{\Theta}_{D}}$ of $\boldsymbol{\Theta}^{*}_{D}$ with a rate $\max_{i=1,\dotsc,p}|\widehat{\boldsymbol{\Theta}_{D}}_{,ii}-\boldsymbol{\Theta}^{*}_{D,ii}|=O_{P}(\sqrt{\log n/n})$. Similarly, SPACE also requires a separate estimator of $\boldsymbol{\Theta}_{D}^{*}$ with the same rate. Such a separate estimator is difficult to find in practice. In this context, the precise connection between CONCORD parameter estimates and the partial correlation remains ambiguous. Computationally, CONCORD-ISTA algorithm [29] possesses a sublinear $O(1/t)$ rate of convergence, where $t$ is the number of iterations and the convergence is measured in terms of the objective function value. In addition to the sublinear convergence of the objective function, the variable iterate $\{\boldsymbol{\Theta}^{(t)}\}$ may converge to the minimizer $\hat{\boldsymbol{\Theta}}$ of (2.2) in an arbitrarily slow rate [2]. Since the algorithm has to be terminated within a finite number of iterations, statistical error of $\boldsymbol{\Theta}^{(t)}$ (say $\|\boldsymbol{\Theta}^{(t)}-\boldsymbol{\Theta}^{*}\|$) may remain quite large compared to that of $\hat{\boldsymbol{\Theta}}$ (say $\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{*}\|)$ even after a large number of iterations.
In the remainder of this paper, we propose in detail a novel approach to address these challenges and illustrate its usefulness through extensive numerical experiments and an analysis of high-dimensional multi-omics data.